OpenStudy (anonymous):

if m angle1=(4x+12), angle2= (7x+3), and m angle 3= 6y then x =? m angle 1=?, and m angle 2 =? and y =?? help!!!

OpenStudy (anonymous):

drawn a picture if you can

OpenStudy (anonymous):



OpenStudy (anonymous):

Angles 1 and 2 will add up to 180 degrees if that is a straight line you drew. So, 4x+12+7x+3=180 11x+15=180 11x = 165 x = 15

OpenStudy (anonymous):

Then angles 2 and 3 will add up to 180 degrees. So, first find out what angle two is by substituting 15 in for x. 7x+3=7(15)+3=105+3=108 Angle 2 + angle 3 = 180 108 + angle 3 = 180, so angle 3 must be 72 degrees

OpenStudy (anonymous):

Angle 3 is 6y 6y = 72 y=72/6 y=12

OpenStudy (anonymous):

ok thank you very much